# Risk Bounds for Mixture Density Estimation on Compact Domains via the $h$-Lifted Kullback–Leibler Divergence

### v1

@@ 2.2 Advantages of the $h$-lifted KL divergence
For probability density functions $f,g$, and $h$, where $h$ is such that $h(x)\geq\gamma>0$ for all $x\in\mathcal{X}$, the following inequality holds:
$\mathrm{KL}_{h}\left(f\,||\,g\right)\leq\gamma^{-1}L_{2}^{2}(f,g).$
Proof.  Defining $\tilde{f}$ and $\tilde{g}$ as above, we have
The following section is devoted to establishing some technical definitions and instrumental results which will be utilized in later sections.
3 Preliminaries
$\mathcal{V}=\text{cl}\left(\bigcup_{k\in\mathbb{N}}\left\{\sum_{j=1}^{k}\varpi_{j}\varphi\left(\cdot;\theta_{j}\right)\mid\varphi\left(\cdot;\theta_{j}\right)\in\mathcal{P},\varpi_{j}\in\mathbb{R},j\in\left[k\right]\right\}\right),$
Lemma 4.
then, for each $k\in\mathbb{N}$,
1:  Input: $\bar{f}_{0}\in\mathcal{P}$
2:  for $k\in\mathbb{N}$ do
5:  end for
We are interested in two choices for $\kappa$:
$\kappa\left(p\right)=\mathrm{KL}_{h}\left(f\,||\,p\right)$ (7)
and its sample counterpart,
where $(X_{i})_{i\in[n]}$ and $(Y_{i})_{i\in[n]}$ are realisations of $X$ and $Y$, respectively. For choice (7), by the dominated convergence theorem, we observe that
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)$
$\qquad\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f+h}{\left(1-\pi\right)p+\pi q+h}\right\}$
$\qquad\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\left\{\frac{\left(p-q\right)^{2}}{\left[\left(1-\pi\right)p+\pi q+h\right]^{2}}\right\}.$
for every $p,q\in\mathcal{C}$ and $\pi\in\left(0,1\right)$, and thus
Similarly, for case (8), we have
$\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa_{n}\left(\left(1-\pi\right)p+\pi q\right)=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f\left(x_{i}\right)+h\left(x_{i}\right)}{\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)}\\ +\frac{1}{n}\sum_{i=1}^{n}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\log\frac{f\left(y_{i}\right)+h\left(y_{i}\right)}{\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)}\\ =\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(x_{i}\right)-q\left(x_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(x_{i}\right)+\pi q\left(x_{i}\right)+h\left(x_{i}\right)\right]^{2}}+\frac{1}{n}\sum_{i=1}^{n}\frac{\left(p\left(y_{i}\right)-q\left(y_{i}\right)\right)^{2}}{\left[\left(1-\pi\right)p\left(y_{i}\right)+\pi q\left(y_{i}\right)+h\left(y_{i}\right)\right]^{2}}.$
$\sup_{p,q\in\mathcal{C},\pi\in\left(0,1\right)}\frac{\mathrm{d}^{2}}{\mathrm{d}\pi^{2}}\kappa\left(\left(1-\pi\right)p+\pi q\right)\leq\frac{2c^{2}}{a^{2}}<\infty,$
as required. We therefore obtain the following result.
Proposition 5.
Let $\square$ denote either $\kappa$, the $\textrm{KL}_{h}$ divergence (7), or $\kappa_{n}$, the sample $\textrm{KL}_{h}$ divergence (8), and assume that $h\geq a$ and $\varphi\left(\cdot;\theta\right)\leq c$, for each $\theta\in\Theta$. Then,
$\square\left(\bar{f}_{k}\right)-\inf_{p\in\mathcal{C}}\square\left(p\right)\leq\frac{4a^{-2}c^{2}}{k+2},$
Notice that sequences $\left(\bar{f}_{k}\right)_{k\in\mathbb{N}}$ obtained via Lemma 4 are greedy approximation sequences, and that $\bar{f}_{k}\in\mathcal{C}_{k}$, for each $k\in\mathbb{N}$. Let $\left(f_{k}\right)_{k\in\mathbb{N}}$ be the sequence of minimizers defined by
Proposition 6.
For the $\textrm{KL}_{h}$ divergence (7), under the assumption that $h\geq a$ and $\varphi\left(\cdot;\theta\right)\leq c$, for each $\theta\in\Theta$, we have
for each $k\in\mathbb{N}$, where $\left(f_{k}\right)_{k\in\mathbb{N}}$ is the sequence of minimizers defined via (9). Furthermore, for the sample $\textrm{KL}_{h}$ divergence (8), under the same assumptions as above, we have
As is common in many statistical learning/uniform convergence results (e.g., Bartlett and Mendelson, 2002, Koltchinskii and Panchenko, 2004), we employ the use of Rademacher processes and associated bounds. Let $(\varepsilon_{i})_{i\in[n]}$ be i.i.d. Rademacher random variables, that is $\mathrm{\mathbf{P}}(\varepsilon_{i}=-1)=\mathrm{\mathbf{P}}(\varepsilon_{i}=1)=\nicefrac{{1}}{{2}}$, that are independent of $(X_{i})_{i\in[n]}$. The Rademacher process, indexed by a class of real measurable functions $\mathcal{S}$, is defined as the quantity
$R_{n}(s)=\frac{1}{n}\sum_{i=1}^{n}s(X_{i})\varepsilon_{i},$
Lemma 7 (Rockafellar, 1997, Thm. 32.2).
In particular, we use the fact that since a linear functional of convex combinations achieves its maximum value at vertices, the Rademacher complexity of $\mathcal{S}$ is equal to the Rademacher complexity of $\textrm{co}(\mathcal{S})$ (see Lemma 19). We consequently obtain the following result.
Lemma 8.
Proof.  Follows immediately from Lemma 7.
@@ 4 Main results
Here we provide explicit statements regarding the convergence rates claimed in (5). We assume that $f$ is bounded above by some constant $c$ and that the lifting function $h$ is bounded above and below by constants $a$ and $b$, respectively.
@@
5 Proofs
We first present a result establishing a uniform concentration bound for the $h$-lifted log-likelihood ratios, which is instrumental in the proof of Theorem 9. Our proofs broadly follow the structure of Rakhlin et al. (2005), modified as needed for the use of $\textrm{KL}_{h}$.
Theorem 11.
Let $X_{1},\ldots,X_{n}$ be an i.i.d. sample of size $n$ drawn from a fixed density $f$ such that $0\leq f(x)\leq c$ for all $x\in\mathcal{X}$, and let $h$ be a positive density with $0<a\leq h(x)\leq b$ for all $x\in\mathcal{X}$. Then with probability at least $1-\mathrm{e}^{-t}$,
@@ 5 Proofs
The bound on the term
Proof. [of Theorem 11]. Fix $h$ and define the following quantities: $\tilde{g}=g+h$, $\tilde{f}=f+h$, $\tilde{\mathcal{C}}=\mathcal{C}+h$,
$\leq\frac{1}{n}\left(\log\frac{c+b}{a}-\log\frac{a}{c+b}\right)$
$=\frac{1}{n}2\log\frac{c+b}{a}=c_{i}.$
so
$2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$ $=2\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\eta(p_{i})\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$
$\leq\frac{4(c+b)}{a}\operatorname*{{\mathrm{\mathbf{E}}}}\left\lVert\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}\varepsilon_{i}-\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\right\rVert_{\tilde{\mathcal{C}}}$
By again applying Lemma 20, we have
Applying Lemmas 8 and 23, the following inequality holds for some constant $K>0$:
$\left\lVert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rVert$
$\qquad\leq\frac{8(c+b)K}{a^{2}\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon+\frac{(8b+4a)(c+b)}{a^{2}\sqrt{n}}+\sqrt{2}\log\left(\frac{c+b}{a}\right)\sqrt{\frac{t}{n}},$
$\qquad=\frac{w_{1}}{\sqrt{n}}\operatorname*{{\mathrm{\mathbf{E}}}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}},$
where $w_{1}$, $w_{2}$, and $w_{3}$ are constants that each depend on some or all of $a$, $b$, and $c$.
Remark 13.
Proposition 14.
By combining Lemmas 16 and 17, the following inequality holds:
$\log N(\mathcal{P},\varepsilon,d_{n,x})\leq\log N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}),$
and
$\log N(\mathcal{P},\varepsilon,d_{n,y})\leq\log N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty}).$
With this result, we can now prove Theorem 9.
Proof. [of Theorem 9] The notation is the same as in the proof of Theorem 11. The values of the constants may change from line to line.
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)-\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)$
$\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}$
$\quad={\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}$
$\quad\qquad+{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\quad\qquad-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}$
$\quad\qquad-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}$
$\quad=\left({\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}\right)+\left({\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k,n}}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}\right)$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{f}}{\tilde{f}_{k}}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{f}}{\tilde{f}_{k}}\right)$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}\right)$
$\quad\leq 2\sup_{\tilde{g}\in\tilde{\mathcal{C}}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(X_{i})}{\tilde{f}(X_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{f}\log\frac{\tilde{g}}{\tilde{f}}\right\rvert+2\sup_{\tilde{g}\in\tilde{\mathcal{C}}}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{g}(Y_{i})}{\tilde{f}(Y_{i})}-{\operatorname*{{\mathrm{\mathbf{E}}}}}_{h}\log\frac{\tilde{g}}{\tilde{f}}\right\rvert$
$\quad\quad+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k,n}(X_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(X_{i})}{\tilde{f}_{k}(X_{i})}\right)+\left(\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}(Y_{i})}{\tilde{f}_{k}(Y_{i})}\right)$
$\quad\leq 2\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\frac{w^{x}_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,x})\mathrm{d}\varepsilon\right\}+\frac{w^{x}_{2}}{\sqrt{n}}+w^{x}_{3}\sqrt{\frac{t}{n}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}$
$\quad\qquad+2\operatorname*{{\mathrm{\mathbf{E}}}}\left\{\frac{w^{y}_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon,d_{n,y})\mathrm{d}\varepsilon\right\}+\frac{w^{y}_{2}}{\sqrt{n}}+w^{y}_{3}\sqrt{\frac{t}{n}}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\quad\leq\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}$
$\quad\qquad+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})},$
$\mathrm{KL}_{h}\left(f_{k}\,||\,f_{k,n}\right)$ $=\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(X_{i})}{\tilde{f}_{k,n}(X_{i})}+\frac{1}{n}\sum_{i=1}^{n}\log\frac{\tilde{f}_{k}(Y_{i})}{\tilde{f}_{k,n}(Y_{i})}$
$\leq\frac{4a^{-2}c^{2}}{k+2}+\inf_{p\in\mathcal{C}}\mathrm{KL}_{h}\left(f_{k}\,||\,p\right).$
$\mathrm{KL}_{h}\left(f\,||\,f_{k,n}\right)$ $-\mathrm{KL}_{h}\left(f\,||\,f_{k}\right)$
$\quad\leq\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}+\frac{w_{4}}{k+2}.$ (14)
$\leq\frac{w_{4}}{k+2}+\frac{w_{1}}{\sqrt{n}}\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon+\frac{w_{2}}{\sqrt{n}}+w_{3}\sqrt{\frac{t}{n}}.$
where $c_{1}$, $c_{2}$, and $c_{3}$ are constants that depend on some or all of $a$, $b$, and $c$.
Remark 15.
Proof. [of Corollary 10] Let $\mathcal{X}$ and $\Theta$ be compact and assume the Lipshitz condition given in (12). If $\varphi(x;\cdot)$ is continuously differentiable, then
$\left|\varphi\left(x;\theta\right)-\varphi\left(x;\tau\right)\right|$ $\leq\sum_{k=1}^{d}\left|\frac{\partial\varphi\left(x;\cdot\right)}{\partial\theta_{k}}\left(\theta_{k}^{*}\right)\right|\left|\theta_{k}-\tau_{k}\right|$
$\leq\sup_{\theta^{*}\in\Theta}\left\|\frac{\partial\varphi\left(x;\cdot\right)}{\partial\theta}\left(\theta^{*}\right)\right\|_{1}\left\|\theta-\tau\right\|_{1}.$
Setting
Since $\Theta\subset\mathbb{R}^{d}$, using the fact that a Euclidean set of radius $r$ has covering number
$N\left(r,\varepsilon\right)\leq\left(\frac{3r}{\varepsilon}\right)^{d},$
we have
So
and since $c<\infty$, the integral is finite, as required.

### v2

@@ 2.2 Advantages of the $h$-lifted KL divergence
For probability density functions $f,\,g,\,$and $h$, where $h$ is such that $h(x)\geq\gamma>0$ for all $x\in\mathcal{X}$, the following inequalities hold:
$\frac{1}{4}L_{1}^{2}\left(f,g\right)\leq\mathrm{KL}_{h}\left(f\,||\,g\right)\leq\gamma^{-1}L_{2}^{2}\left(f,g\right).$
Proof.
See Appendix C.3. ∎
Remark 4.
Proposition 2 highlights the benefit of the $h$-lifted KL divergence being bounded for all continuous densities, unlike the standard KL divergence, while satisfying a relationship similar to that between the KL divergence and the $L_{2}$ norm distance. Moreover, the first inequality of Proposition 3 is a Pinsker-like relationship between the $h$-lifted KL divergence and the total variation distance $\mathrm{TV}(f,g)=\frac{1}{2}L_{1}(f,g)$.
@@ 3 Main results
Here we provide explicit statements regarding the convergence rates claimed in (6) via Theorem 5 and Corollary 6, which are proved in Appendix A.2. We assume that $f$ is bounded above by some constant $c$ and that the lifting function $h$ is bounded above and below by constants $a$ and $b$, respectively.
@@
Our results are applicable to any compact metric space $\mathcal{X}$, with $\left[0,1\right]$ used in the experimental setup in Section 4.2 as a simple and tractable example to illustrate key aspects of our theory. There is no issue in generalizing to $\mathcal{X}=\left[-m,m\right]^{d}$ for $m>0$ and $d\in\mathbb{N}$, or more abstractly, to any compact subset $\mathcal{X}\subset\mathbb{R}^{d}$. Additionally, $\mathcal{X}$ could even be taken as a functional compact space, though establishing compactness and constructing appropriate component classes $\mathcal{P}$ over such spaces to achieve small approximation errors $\mathrm{KL}_{h}\left(f\,||\,\mathcal{C}\right)$ is an approximation theoretic task that falls outside the scope of our work.
From the proof of Theorem 5 in Appendix A.2, it is clear that the dimensionality of $\mathcal{X}$ only influences our bound through the complexity of the class $\mathcal{P}$, specifically, the constant $\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon$, which remains independent of both $n$ and $k$. Here, $N(\mathcal{P},\varepsilon,\operatorname*{\lVert\,\cdot\,\rVert})$ is the $\varepsilon$-covering number of $\mathcal{P}$. In fact, the constant with respect to $k$ ($u_{1}$ in Theorem 5) is entirely unaffected by the dimensionality of $\mathcal{X}$. Thus, the rates of our bound on the expected $h$-lifted KL divergence are dimension-independent and hold even when $\mathcal{X}$ is infinite-dimensional, as long as there exists a class $\mathcal{P}$ such that $\int_{0}^{c}\log^{1/2}N(\mathcal{P},\varepsilon/2,{\operatorname*{\lVert\,\cdot\,\rVert}}_{\infty})\mathrm{d}\varepsilon$ is finite.
Corollary 6 provides a method for obtaining such a bound when the elements of $\mathcal{P}$ satisfy a Lipschitz condition.